OpenStudy (anonymous):

du=-3 dx then we get dx =______?

OpenStudy (anonymous):

i though it would be -1/3

OpenStudy (anonymous):

but its not

OpenStudy (anonymous):

It is.

OpenStudy (anonymous):

its not it....i dont get it...

OpenStudy (anonymous):

Or rather it is: \[du = -3 dx \implies dx = -\frac{1}{3}du\]

OpenStudy (anonymous):

Yes, I'm afraid it is.

OpenStudy (anonymous):

ok thanks polpak : )

OpenStudy (anonymous):

polpak i need to complet this step: \[u= 2-3x \int\limits dx/2-3x= \int\limits 1/u (-1/3 du) = (1/3) \int\limits 1/u \]

OpenStudy (anonymous):

du

OpenStudy (anonymous):

I don't understand what you have there...

OpenStudy (anonymous):

Are you trying to solve \[\int \frac{dx}{2-3x}\] ?

OpenStudy (anonymous):

that becomes 1/u...but i need to put it in the final answer as (-1/3) (1/u) + C

OpenStudy (anonymous):

The integral of 1/u du is the natural log of u plus a constant.

OpenStudy (anonymous):

so the natural log of 2 -3x is.....

OpenStudy (anonymous):

That's it.

OpenStudy (anonymous):

Unless you have some limits you can plug in for your x.

OpenStudy (anonymous):

so it would become (-1/3)(2-3x)+C

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

\[-\frac{1}{3}ln(2-3x) + C\]

OpenStudy (anonymous):

there's no limits

OpenStudy (anonymous):

o so then what i do with it? do i take the ln of -1/3?

OpenStudy (anonymous):

No. you don't do anything. that's all there is.

OpenStudy (anonymous):

\[\text{Let } u = 2-3x\]\[\implies du = -3dx\]\[\implies dx = -\frac{1}{3}du\]\[\implies \int \frac{dx}{2-3x} = \int \frac{1}{u}(-\frac{1}{3} du)\]\[ =-\frac{1}{3}ln(2-3x) + C\]

OpenStudy (anonymous):

this is what its asking me to do : -1/3 ln\[\left| u \right|\] + C =_______

OpenStudy (anonymous):

Yes fine, technically the () should be ||

OpenStudy (anonymous):

and i put 2-3x but its wrong

OpenStudy (anonymous):

yes, because -1/3ln|u| + C is not 2-3x.

OpenStudy (anonymous):

Just replace the u with what u equals.

OpenStudy (anonymous):

but you have to write the whole thing.

OpenStudy (anonymous):

2+3x?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

i am doing this online and the system telling me it wrong when i put (-1/3)ln(2-3x)+C....

OpenStudy (anonymous):

well the answer is {-(ln(2-3x))/3}+C

OpenStudy (anonymous):

Those two things are the same.

OpenStudy (anonymous):

So your system is dumb if it can't tell that they are equivalent.

OpenStudy (anonymous):

i know right : (